a student opens a book to two facing pages. The product of the page number is 506. what are the page numbers.
Answer by stanbon(75887)

a student opens a book to two facing pages. The product of the page number is 506. what are the page numbers.
------------
The page numbers are consecutive whole numbers:
x and x+1
------------
Equation:
x * (x+1) = 506
x^2 + x - 506 = 0
(x+23)(x-22) = 0
x = 22
x+1 = 23
==============
Cheers,
Stan H.